OpenStudy (anonymous):

Solve –2(x – 1) = 16

OpenStudy (shaik0124):

-2x+2=16 -2x+2-2=16-2 -2x=14 x=-14/2 x=-7

OpenStudy (anonymous):

Wow, lol idk how yall do it! i wouldnt even think of these complicated steps!